# Constraint-Driven Multi-USV Coverage Path Generation for Aquatic Environmental Monitoring

### v1

@@ I Introduction
Additionally, the kinematic vehicle models assumed in the above papers allow instantaneous velocity changes, but USVs may find problems to follow the velocities generated by coverage controllers due to their greater inertia. Another issue is that water vehicles may suffer significant disturbances due to currents, waves, and wind. For these reasons, in marine craft control it is common to employ a hierarchical architecture composed of control layers for guidance, navigation, and control [30]. Indeed, multiple works follow a backstepping approach [31, 32], where the overall control problem is decoupled into a high-level problem that generates ideal paths to follow and a low-level problem that makes USVs follow the path while rejecting disturbances. That is, the fleet trajectory design is decoupled from the individual trajectory tracking problem, as happens, for example, in [33]. With this in mind, it is not surprising that most results on coverage control with USVs focus on coverage path planning [34, 35, 36], leaving path following to low-level controllers. Nevertheless, even in these simpler setups, there are challenges to solve. For example, it is difficult to update the paths flexibly under real situations because path planning is executed in a different layer. In addition, the path-planning approach generates only a finite-length path, which may not be suitable for persistent monitoring of the aquatic environment.
@@
In this article, we consider an online coverage path generation with real-time feedback on the states of the USVs. Specifically, we deal with circular and elliptic paths as the most basic path shapes. We then formulate a novel online optimal path generation problem corresponding to each path shape, which is based on the formulation of persistent coverage control [14, 15, 16]. To guarantee coverage performance for the path, we transform the specification for minimizing costs into a constraint-based specification. It is then pointed out that the problem may suffer from a combinatorial explosion associated with binary variables corresponding to the turning directions of the USVs. To address this issue, we present a more conservative scheme to reduce computational complexity, which is also shown to provide a partially distributed structure. We then present a partially distributed constraint-based control as in [15, 21, 22, 23] to meet the performance constraint together with constraints that parameters of the path are inherently required to satisfy. The proposed online coverage path generator is then demonstrated with an ideal mathematical model over a wide area free from the space constraints. We finally implement a hierarchical control architecture including the present coverage path generator on an experimental testbed, and demonstrate that the path generator works even in the presence of various uncertainties in the real physical world.
@@
    To the best of the authors’ knowledge, this article is the first to apply the concept of constraint-based control to the high-level path generation. Indeed, the constraint-based control is generally used as low-level safe control.
@@ II Problem Formulation
In this article, we design an online path generator to efficiently sample data on all points $q_{j}\ (j=1,2,\dots,m)$, assuming two types of closed paths, namely circular and elliptic paths, so as to reduce the number of parameters that characterize the path.
@@ II-A Circular Path Generation Problem
respectively. Solving the quadratic equation (5a) together with (5b) for $c_{i}$ yields two solutions, which correspond to the right-hand turning path, denoted by $\mathrm{r}$, and left-hand one, denoted by $\mathrm{l}$, as illustrated in Fig. 2. Thus, the turning direction becomes another decision variable for each USV $i$, which is denoted by ${\rm X}_{i}\in\mathcal{X}_{i}=\{\mathrm{r},\mathrm{l}\}$. Notice that once the turning direction ${\rm X}_{i}\in\mathcal{X}_{i}$ is fixed, we can eliminate the variable $c_{i}$ by solving (5) in advance. Specifically, given $z_{i}(t)=[p_{i}^{\top}(t)\ \theta_{i}(t)]^{\top}$, the center $c_{i}$ meeting (5) is given by
@@
Figure 2: Two possible circular paths with right-hand and left-hand turning.
@@
Accordingly, the value $f(p^{*}_{ij}(c_{i},r_{i}),q_{j})$ provides a measure on the quality of the path ${\mathcal{P}}_{\rm c}(c_{i},r_{i})$. Nevertheless, the path is changing in real time, so the USV may not actually visit point $p^{*}_{ij}(c_{i},r_{i})$. This motivates us to scale $f(p^{*}_{ij}(c_{i},r_{i}),q_{j})$ depending on the time required for the USV at $p_{i}$ to arrive at $p^{*}_{ij}(c_{i},r_{i})$, which is proportinal to the arc angle from $p_{i}$ to $p^{*}_{ij}(c_{i},r_{i})$, defined as
@@
where $\mathrm{X}_{i}\in\mathcal{X}_{i}$.
@@ II-B Elliptic Path Generation Problem
respectively. Analogous to the analysis in Section II-A, the solution to (15) for $c_{i}$ yields two alternatives, each corresponding to a turning direction $\mathrm{X}_{i}\in\mathcal{X}_{i}$, as depicted in Fig. 3. Upon fixing the direction $\mathrm{X}_{i}$, the center $c_{i}$ satisfying (15) is uniquely determined as a function of $s_{i}$. Specifically, when $\mathrm{X}_{i}=\mathrm{r}$, the center $c_{i}$ meeting (15) for a given $z_{i}(t)$ is formulated as
@@ III-A Circular Coverage Path Generator
The goal of this section is thus to determine a pair of $r$ and $\mathrm{X}\in\mathcal{X}$ so that the inequality
@@
is satisfied for all time $t\geq 0$, where $\mathcal{X}=\mathcal{X}_{1}\times\cdots\times\mathcal{X}_{n}$.
@@
Now, the constraint (26) essentially comprises $2^{n}$ combinations corresponding to the elements of the set $\mathcal{X}$, which may cause computational problems. We thus consider an alternative constraint while accepting conservatism. To this end, let us define a Voronoi-like partition of the finite set $\mathcal{Q}$ as
@@
$J^{\mathrm{X}}_{\mathrm{c}}(r;z(t),\phi(t))=\sum_{j:q_{j}\in\mathcal{V}_{\mathrm{c},i}(r,\mathrm{X};z(t))}g^{{\rm X}_{i}}_{{\rm c},ij}(r_{i};z_{i}(t))\phi_{j}(t).$ (29)
@@
$b^{1}_{\mathrm{c},i}(r_{i};z_{i}(t),\phi(t))=\max_{{\rm X}_{i}\in\mathcal{X}_{i}}I^{{\rm X}_{i}}_{\mathrm{c},i}(r_{i};z_{i}(t),\phi(t))-\frac{\gamma}{n},$ (30)
@@
Suppose now that $r_{i}$ is selected so that $b^{1}_{\mathrm{c},i}(r_{i};z_{i}(t),\phi(t))\geq 0$ holds, and that $\mathrm{X}_{i}\in\mathcal{X}_{i}$ is determined by
$\mathrm{X}_{i}=\arg\max_{{\rm X}_{i}\in\mathcal{X}_{i}}I^{{\rm X}_{i}}_{\mathrm{c},i}(r_{i};z_{i}(t),\phi(t))$ (32)
@@
For any $\mathrm{X}\in\mathcal{X}$, we have that
@@
The function $b^{1}_{\mathrm{c},i}$ in (30) includes only $\mathrm{X}_{i}$, which unlike (26) has two options, $\mathrm{r}$ and $\mathrm{l}$, drastically reducing the computational effort for meeting the associated inequality constraints. Consequentially, we take $b^{1}_{\mathrm{c},i}\geq 0\ (i=1,2,\dots,n)$ instead of (26) at the cost of conservatism. Besides the reduction of the computational effort, the constraint $b^{1}_{\mathrm{c},i}\geq 0\ (i=1,2,\dots,n)$ allows the partially distributed implementation of the present algorithm. That is, given $\mathcal{V}_{\mathrm{c},i}^{-}(t)$ from a central computer, the constraint $b^{1}_{\mathrm{c},i}\geq 0$ is handled locally by USV $i\ (i=1,2,\dots,n)$, and (32) is also locally executed by USV $i$. The overall control architecture is illustrated in Fig. 4.
@@ III-B Elliptic Coverage Path Generator
$J^{\mathrm{X}}_{\mathrm{e}}(s;z(t),\phi(t))=\sum_{j:q_{j}\in\mathcal{V}_{\mathrm{e},i}(s,\mathrm{X};z(t))}g^{{\rm X}_{i}}_{\mathrm{e},ij}(s_{i};z_{i}(t))\phi_{j}(t)$ (42)
@@
$b^{1}_{\mathrm{e},i}(s_{i};z_{i}(t),\phi(t))=\max_{{\rm X}_{i}\in\mathcal{X}_{i}}I^{{\rm X}_{i}}_{\mathrm{e},i}(s_{i};z_{i}(t),\phi(t))-\frac{\gamma}{n},$ (43)
@@
Suppose now that $s_{i}$ is selected so that $b^{1}_{\mathrm{e},i}(s_{i};z_{i}(t),\phi(t))\geq 0$ holds, and that $\mathrm{X}_{i}\in\mathcal{X}_{i}$ is determined by
$\mathrm{X}_{i}=\arg\max_{{\rm X}_{i}\in\mathcal{X}_{i}}I^{{\rm X}_{i}}_{\mathrm{e},i}(s_{i};z_{i}(t),\phi(t))$ (45)
@@
where $w_{i}$ is a slack variable, $\lambda>0$ is a penalty parameter for violations on $b^{1}_{\mathrm{e},i}\geq 0$, and $\alpha_{1}$ and $\alpha_{2,\dots,5}$ are extended class $\mathcal{K}$ functions. The quadratic programming representation corresponding to (46) is shown in Appendix B. Once $s_{i}$ is updated by the designed $\rho_{i}$, we then update ${\rm X}_{i}$ based on (45). When the variables $s_{i}$ and ${\rm X}_{i}$ are updated, the center $c_{i}$ is computed using either (16) or (17), depending on whether ${\rm X}_{i}=\mathrm{r}$ or ${\rm X}_{i}=\mathrm{l}$.
@@
$\omega_{i}^{*}$ $=\zeta({\mathrm{X}}_{i})\bar{v}\kappa(p_{i},c_{i},S_{i}),$ (47)
$\kappa(p_{i},c_{i},S_{i})$ $=\frac{|r_{x}r_{y}|}{r_{x}^{2}\sin^{2}{\varphi_{S}}+r_{y}^{2}\cos^{2}{\varphi_{S}}},$ (48)
@@
where $\kappa(p_{i},c_{i},S_{i})$ represents the curvature of the elliptic path ${\mathcal{P}}_{\rm e}(c_{i},S_{i})$ at $p_{i}$, $r_{x}$ and $r_{y}$ are the reciprocal number of the eigenvalues of $S_{i}$, and $\vartheta_{S}$ is a rotation angle of an eigenvector $v_{r_{x}}$, corresponding to an eigenvalue $r_{x}$, which is given as $\vartheta_{S}=\operatorname*{atan2}2(\mathbf{e}_{2}^{\top}v_{r_{x}},\mathbf{e}_{1}^{\top}v_{r_{x}})$.

### v2

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Additionally, the kinematic vehicle models assumed in the above papers allow instantaneous velocity changes, but USVs may find problems to follow the velocities generated by coverage controllers due to their greater inertia. Another issue is that water vehicles may suffer significant disturbances due to currents, waves, and wind. For these reasons, in marine craft control it is common to employ a hierarchical architecture composed of control layers for guidance, navigation, and control [30]. Indeed, multiple works follow a backstepping approach [31, 32], where the overall control problem is decoupled into a high-level problem that generates ideal paths to follow and a low-level problem that makes USVs follow the path while rejecting disturbances. That is, the fleet trajectory design is decoupled from the individual trajectory tracking problem, as happens, for example, in [33]. With this in mind, it is not surprising that most results on coverage control with USVs focus on coverage path planning [34, 35, 36, 37], leaving path following to low-level controllers. Nevertheless, even in these simpler setups, there are challenges to solve. For example, it is difficult to update the paths flexibly under real situations because path planning is executed in a different layer. In addition, the path-planning approach generates only a finite-length path, which may not be suitable for persistent monitoring of the aquatic environment. Note that the concept of the constraint-based control has been recently applied to path planning or trajectory generation e.g. in [38, 39, 40, 41, 42]. However, the above shortcomings of [34, 35, 36, 37] apply to the papers [38, 39, 40, 41], and all of [38, 39, 41, 42] focus on safety certificates and cannot be a solution to the area coverage problem.
@@
In this article, we consider an online coverage path generation with real-time feedback on the states of the USVs. Specifically, we deal with circular and elliptic paths as the most basic path shapes. We then formulate a novel online optimal path generation problem corresponding to each path shape, which is based on the formulation of persistent coverage control [14, 15, 16]. To guarantee coverage performance for the path, we transform the specification for minimizing costs into a constraint-based specification. It is then pointed out that the problem may suffer from a combinatorial explosion associated with binary variables corresponding to the turning directions of the USVs. To address this issue, we present a more conservative scheme to reduce computational complexity, which is also shown to provide a partially distributed structure. We then present a partially distributed constraint-based control as in [15, 21, 22, 23] to meet the performance constraint together with constraints that parameters of the path are inherently required to satisfy. The proposed online coverage path generator is then demonstrated with an ideal mathematical model over a wide area free from the space constraints. We finally implement a hierarchical control architecture including the present coverage path generator on an experimental testbed, and demonstrate that the path generator works even in the presence of various uncertainties in the real physical world. The result is compared with a lawnmower pattern algorithm similar to [37].
@@
    We apply the constraint-based control to the high level path generation, merging path generation and path following layers by exploiting a suitable trajectory parametrization for the nature of the robots considered. This can provide advantages because it combines the best of the two layers: coverage-performance informed decision making and real time feedback.
@@ II Problem Formulation
In this article, we design an online path generator to efficiently sample data on all points $q_{j}\ (j=1,2,\dots,m)$. In principle, our solution can be applied to any path that meet the following conditions:
  1. 1.
    The path should be characterized by a finite number of parameters.
  2. 2.
    Some parameters are successfully eliminated by equality constraints to ensure that the USV can follow the path despite the motion constraints.
  3. 3.
    The point on the path closest to each observation point can be explicitly given as a function of the parameters.
  4. 4.
    The geodesic distance between any two distinct points on the path can be explicitly given as a function of the parameters.
In this paper, we employ the circular and elliptic paths as candidates to meet these constraints due to the following reasons:
  1. 1.
    The number of parameters to specify the path shape is limited, which is advantageous in computational complexity.
  2. 2.
    These parametrization models are well aligned with the system dynamics and easy to implement with the USVs.
  3. 3.
    More complex path shapes make it harder to satisfy the latter two conditions.
Note that the present solution updates the path and hence the actual trajectory that each USV follows can be more complex than the circular or elliptic path.
@@ II-A Circular Path Generation Problem
respectively. Solving the quadratic equation (5a) together with (5b) for $c_{i}$ yields two solutions, which correspond to the right-hand turning path, denoted by $\mathrm{r}$, and left-hand one, denoted by $\mathrm{l}$, as illustrated in Fig. 2. Thus, the turning direction becomes another decision variable for each USV $i$, which is denoted by ${\rm X}_{i}\in\mathcal{X}=\{\mathrm{r},\mathrm{l}\}$. Notice that once the turning direction ${\rm X}_{i}\in\mathcal{X}$ is fixed, we can eliminate the variable $c_{i}$ by solving (5) in advance. Specifically, given $z_{i}(t)=[p_{i}^{\top}(t)\ \theta_{i}(t)]^{\top}$, the center $c_{i}$ meeting (5) is given by
@@
Figure 2: Two possible circular paths with right-hand and left-hand turning. The two red bands indicate important areas.
@@
Accordingly, the value $f(p^{*}_{ij}(c_{i},r_{i}),q_{j})$ provides a measure on the quality of the path ${\mathcal{P}}_{\rm c}(c_{i},r_{i})$. Nevertheless, the path is changing in real time, so the USV may not actually visit point $p^{*}_{ij}(c_{i},r_{i})$. For example, suppose that a USV is facing a decision on the better path between $\mathrm{X}_{i}=\mathrm{l}$ or $\mathrm{X}_{i}=\mathrm{r}$, as shown in Fig. 2, where both of the regions A and B are assumed to have high importance indices. Then, the path with $\mathrm{X}_{i}=\mathrm{l}$ looks better because it would drive the USV to the important area more quickly. Also, notice that the alternative choice $\mathrm{X}_{i}={\rm r}$ might have been updated before arriving at region B, so that the path $\mathrm{X}_{i}={\rm l}$ should be prioritized. This motivates us to scale $f(p^{*}_{ij}(c_{i},r_{i}),q_{j})$ depending on the time required for the USV at $p_{i}$ to arrive at $p^{*}_{ij}(c_{i},r_{i})$, which is proportional to the arc angle from $p_{i}$ to $p^{*}_{ij}(c_{i},r_{i})$, defined as
@@
where $\mathrm{X}_{i}\in\mathcal{X}$.
@@ II-B Elliptic Path Generation Problem
respectively. Analogous to the analysis in Section II-A, the solution to (15) for $c_{i}$ yields two alternatives, each corresponding to a turning direction $\mathrm{X}_{i}\in\mathcal{X}$, as depicted in Fig. 3. Upon fixing the direction $\mathrm{X}_{i}$, the center $c_{i}$ satisfying (15) is uniquely determined as a function of $s_{i}$. Specifically, when $\mathrm{X}_{i}=\mathrm{r}$, the center $c_{i}$ meeting (15) for a given $z_{i}(t)$ is formulated as
@@ III-A Circular Coverage Path Generator
The goal of this section is thus to determine a pair of $r$ and $\mathrm{X}\in\bar{\mathcal{X}}$ so that the inequality
@@
is satisfied for all time $t\geq 0$, where $\bar{\mathcal{X}=\mathcal{X}^{n}}$.
@@
Now, the constraint (26) essentially comprises $2^{n}$ combinations corresponding to the elements of the set $\bar{\mathcal{X}}$, which may cause computational problems. We thus consider an alternative constraint while accepting conservatism. To this end, let us define a Voronoi-like partition of the finite set $\mathcal{Q}$ as
@@
$J^{\mathrm{X}}_{\mathrm{c}}(r;z(t),\phi(t))=\sum_{i=1}^{n}\sum_{j:q_{j}\in\mathcal{V}_{\mathrm{c},i}(r,\mathrm{X};z(t))}g^{{\rm X}_{i}}_{{\rm c},ij}(r_{i};z_{i}(t))\phi_{j}(t).$ (29)
@@
$b^{1}_{\mathrm{c},i}(r_{i};z_{i}(t),\phi(t))=\max_{{\rm X}_{i}\in\mathcal{X}}I^{{\rm X}_{i}}_{\mathrm{c},i}(r_{i};z_{i}(t),\phi(t))-\frac{\gamma}{n},$ (30)
@@
Suppose now that $r_{i}$ is selected so that $b^{1}_{\mathrm{c},i}(r_{i};z_{i}(t),\phi(t))\geq 0$ holds, and that $\mathrm{X}_{i}\in\mathcal{X}$ is determined by
$\mathrm{X}_{i}=\arg\max_{{\rm X}_{i}\in\mathcal{X}}I^{{\rm X}_{i}}_{\mathrm{c},i}(r_{i};z_{i}(t),\phi(t))$ (32)
@@
For any $\mathrm{X}\in\bar{\mathcal{X}}$, we have that
@@
In order to directly evaluate the original function $J^{\rm X}_{\rm c}(r;z(t),\phi(t))$ in (29) in computing $\rho_{i}\ (i=1,2,\dots,n)$, we in principle have to consider all variations of $\mathcal{V}_{{\rm c},i}(r,\mathrm{X};z(t))$ corresponding to every $\mathrm{X}\in\bar{\mathcal{X}}$, which is computationally expensive or even unrealistic. On the other hand, the functions (30), (31) are free from such variations, since the partition is fixed to $\mathcal{V}^{-}_{{\rm c},i}$. Needless to say, there is a gap between the Voronoi-like partition for $\mathrm{X}_{i}$ updated by (32) and the collection of $\mathcal{V}^{-}_{{\rm c},i}$. Theorem 1 ensures that the original constraint (26) is satisfied by evaluating $b^{1}_{{\rm c},i}(r;z(t),\phi(t))$, even in the presence of the gap in the partition. Now, the definition of $b^{1}_{\mathrm{c},i}$ in (30) includes only $\mathrm{X}_{i}$, which unlike (26) has two options, $\mathrm{r}$ and $\mathrm{l}$, drastically reducing the computational effort for meeting the associated inequality constraints. Consequentially, we take $b^{1}_{\mathrm{c},i}\geq 0\ (i=1,2,\dots,n)$ instead of (26) at the cost of conservatism. Besides the reduction of the computational effort, the constraint $b^{1}_{\mathrm{c},i}\geq 0\ (i=1,2,\dots,n)$ allows the partially distributed implementation of the present algorithm [^3] There is no clear consensus on how to call the present architecture, but we take partially distributed following [44, 45, 46].. That is, given $\mathcal{V}_{\mathrm{c},i}^{-}(t)$ from a central computer, the constraint $b^{1}_{\mathrm{c},i}\geq 0$ is handled locally by USV $i\ (i=1,2,\dots,n)$, and (32) is also locally executed by USV $i$. The overall control architecture is illustrated in Fig. 4.
@@ III-B Elliptic Coverage Path Generator
$J^{\mathrm{X}}_{\mathrm{e}}(s;z(t),\phi(t))=\sum_{i=1}^{n}\sum_{j:q_{j}\in\mathcal{V}_{\mathrm{e},i}(s,\mathrm{X};z(t))}g^{{\rm X}_{i}}_{\mathrm{e},ij}(s_{i};z_{i}(t))\phi_{j}(t)$ (42)
@@
$b^{1}_{\mathrm{e},i}(s_{i};z_{i}(t),\phi(t))=\max_{{\rm X}_{i}\in\mathcal{X}}I^{{\rm X}_{i}}_{\mathrm{e},i}(s_{i};z_{i}(t),\phi(t))-\frac{\gamma}{n},$ (43)
@@
Suppose now that $s_{i}$ is selected so that $b^{1}_{\mathrm{e},i}(s_{i};z_{i}(t),\phi(t))\geq 0$ holds, and that $\mathrm{X}_{i}\in\mathcal{X}$ is determined by
$\mathrm{X}_{i}=\arg\max_{{\rm X}_{i}\in\mathcal{X}}I^{{\rm X}_{i}}_{\mathrm{e},i}(s_{i};z_{i}(t),\phi(t))$ (45)
@@
where $w_{i}$ is a slack variable, $\lambda>0$ is a penalty parameter for violations on $b^{1}_{\mathrm{e},i}\geq 0$, and $\alpha_{1}$ and $\alpha_{2,\dots,5}$ are extended class $\mathcal{K}$ functions. Note that it is not difficult to confirm that there is always $\rho_{i}$ meeting all of the constraints for $k=2,3,4,5$ and hence they are treated as hard constraints. They may conflict only with the constraint with $k=1$ and we relax only this constraint as a soft constraint. The quadratic programming representation corresponding to (46) is shown in Appendix B. Once $s_{i}$ is updated by the designed $\rho_{i}$, we then update ${\rm X}_{i}$ based on (45). When the variables $s_{i}$ and ${\rm X}_{i}$ are updated, the center $c_{i}$ is computed using either (16) or (17), depending on whether ${\rm X}_{i}=\mathrm{r}$ or ${\rm X}_{i}=\mathrm{l}$.
@@
$\omega_{i}^{*}$ $=\zeta({\mathrm{X}}_{i})\bar{v}\kappa_{i}(p_{i},c_{i},S_{i}),$ (47)
$\kappa_{i}(p_{i},c_{i},S_{i})$ $=\frac{|\varrho_{x,i}\varrho_{y,i}|}{{\varrho_{x,i}}^{2}\sin^{2}{\varphi_{S_{i}}+\varrho_{y,i}}^{2}\cos^{2}{\varphi_{S_{i}}}},$ (48)
@@
where $\kappa_{i}(p_{i},c_{i},S_{i})$ represents the curvature of the elliptic path ${\mathcal{P}}_{\rm e}(c_{i},S_{i})$ at $p_{i}$, $\varrho_{x,i}$ and $\varrho_{y,i}$ are the reciprocal number of the eigenvalues of $S_{i}$, and $\vartheta_{S_{i}}$ is a rotation angle of an eigenvector $v_{\varrho_{x,i}}$, corresponding to an eigenvalue $\varrho_{x,i}$, which is given as $\vartheta_{S_{i}}=\operatorname*{atan2}2(\mathbf{e}_{2}^{\top}{v_{\varrho_{x,i}}},\mathbf{e}_{1}^{\top}v_{\varrho_{x,i}})$.
Remark 1
Especially for large USV fleets, no observation point $q_{j}$ may be assigned to a USV as a result of partitioning in (28) or (41). When this happens, all the coefficients for the variable $\rho_{i}$ in the inequality constraint on the performance guarantee in (33) or (46) get equal to zero, and hence the constraint gets independent of the selection of $\rho_{i}$. Noticing that the cost in the QP is given by $\|\rho_{i}\|^{2}$ plus the cost for the slack variable, the optimal solution must be $\rho_{i}=0$ and hence the previous path parameters must be maintained. Thus, even when no observation point is assigned to a USV, no serious problem happens except for the violation of the local performance constraint.
Remark 2
The Voronoi-like partition may have instantaneous big changes and accordingly the paths may significantly change, especially for large USV fleets. Even in the presence of such changes, the paths must continue to meet (5) or (15) as well as the size constraints (27) or (40). Consequently, such changes do not impose extreme actions on the USVs. In other words, the effect of possible instability of the Voronoi-based allocation is inherently mitigated by these constraints.